OpenStudy (anonymous):

find the limit of the function limit sin3x/8x x->0

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

No problemo! Have fun! use l'hopital sin3x turns into 3(cosx) and 8x derives into 8 so, the answer is 3/8 Do let me know if you don't understand it. I'd love to explain it since I just learned this this morning. Hee hee.

OpenStudy (anonymous):

i wanted to ask how does the sin3x turn into 3(cosx)?

OpenStudy (anonymous):

Sure thing. The derivative of sinx is always cosx, right? (It's one of the 'formulas' you should memorize). The 3 in front of X may be moved in front of the sin, so you have 3sinX. After using l'hopital's rule, you get 3(cosx).

OpenStudy (anonymous):

The derivative of sin(3x) is NOT 3cos(x) its 3cos(3x).

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

\[\frac{3}{8}\] with your eyeballs

OpenStudy (anonymous):

you can solve this using either L hospital's rule or in simple way:- to apply L hospital rule it is necessary to check that function should give 0/0or (infinity/infinity) when we give x=0 here when we put x=0 in sin3x/8x it gives 0/0 hence we can apply L hospital's rule; what rule says is just differentiate numerator and denominator until you get the constant value; -> (d/dx (sin3x) ) / (d/dx of 8x) -> 3cos3x /8 ...now see your denominator is free from x i.e if you put x=0 now your denominator can't give 0; put x=0 above -> 3 cos0 /8 =3/8 answer

OpenStudy (anonymous):

you need no l'hopital or anything of that sort. you can see it clearly, so forget l'hopital

OpenStudy (anonymous):

lhopital is so much easier than anything else, though!

OpenStudy (anonymous):

only one thing you need to know, namely that \[lim_{x\rightarrow 0} \frac{\sin(x)}{x}=1\]

OpenStudy (anonymous):

@uber yes but that is like saying that it is so much easier to kill a mouse with an M16

OpenStudy (anonymous):

if you have already learned lots about derivatives etc. then you can apply l'hopital. but typically these questions arrive in a calc text before even derivatives. so there is no using l'hopital.

OpenStudy (anonymous):

@galactic what are you presumed to know before you can answer this question?

OpenStudy (anonymous):

that sin(x) = -cos(x) that i remember we learned in class the other day

OpenStudy (anonymous):

That's not a great example, though. Your example is using a difficult method compared to my uber-easy method. It's more like I'm using a grenade to kill a mouse. Using a M16 is more like using the difference quotient vs using your method. Mine is FARRRR easier than any other method.

OpenStudy (anonymous):

cos(x) = -sin(x) that's what i meant

OpenStudy (anonymous):

Oh phew. Scared me there for a moment. :P Though: You mean to say "the DERIVATIVE of cosx is -sinx" Because cosx doesn't equal -sinx. Make sure you know the difference. ^_^

OpenStudy (anonymous):

hahah yeah i confused myself for a minute....

OpenStudy (anonymous):

if you have not yet taken the derivative of anything, let alone used l'hopital (Hospital) rule you simply do this: you know that \[lim_{x\rightarrow 0}\frac{\sin(x)}{x}=1\] so of course \[lim_{x\rightarrow 0}\frac{\sin(3x)}{3x}=1\] and so \[lim_{x\rightarrow 0}\frac{8\sin(3x)}{3x}=8\] therefore \[lim_{x\rightarrow 0} \frac{\sin(3x)}{8x}=\frac{3}{8}\]

OpenStudy (anonymous):

yea i haven't use the l'hopital yet so the examples you put really help understand

OpenStudy (anonymous):

that is what i assumed. hope the method is both easy and clear.

OpenStudy (anonymous):

yeah it is...the trig functions are sometimes tricky so i take time to understand those one more in depth

OpenStudy (anonymous):

with a small (very small) amount of practice you do it with your eyeballs \[lim_{x\rightarrow 0}\frac{\sin(ax)}{b}=\frac{a}{b}\]

OpenStudy (anonymous):

typo i meant \[lim_{x\rightarrow 0}\frac{\sin(ax)}{bx}=\frac{a}{b}\]

OpenStudy (anonymous):

haha its okay..and yea the book problems help alot so my understanding of this stuff is better though a month of calculus is a bit complex the work keeps me active

OpenStudy (anonymous):

mind wise*